5 times the first of three consecutive even integers is 4 less than twice the sum if the second and third, find the integers
a^5 -4 this is as far as I got

Answer by mananth(12269)

let the first integer be x
the second will be x+2
The third will be x+4
.
5 times the first integer = 5x.
sum of second and third = x+2 + x+4
= 2x+6
Twice the sum of second and third
2(2x+6)
4 less than 2(2x+6)
= 2(2x+6)-4
.
5x=2(2x+6)-4
5x=4x+12-4
x= 8
.
The numbers are 8 , 10, 12